### v1

Abstract
Dirichlet process mixture models (DPMM) provide a principled framework for density estimation and data clustering, offering a robust alternative to parametric mixture models by estimating cluster number $K$ from an infinite choice of partitions induced by a Dirichlet Process $DP(\alpha,G_{0})$ prior on the data. Posterior consistency is guaranteed by an adaptive estimation of the DP concentration parameter $\alpha$, and the estimated $K$ jointly depends on $\alpha$ as well as the variance hyper-parameter of DP base distribution $G_{0}$. Although Markov Chain Monte Carlo (MCMC) methods have bridged the gap between theory and application for such models, they scale poorly to high-dimensional data (e.g. with several hundred or even thousands of features) with limited sample size and suffer from slow and difficult convergence. Variational Inference (VI) exist as an alternative for faster convergence, but lack implementation with increasing model fidelity. Collapsed VI is implemented to integrate out hierarchical parameters beyond mean field VI’s independence assumptions, however, the resulting complexity imposes an implementational barrier to a fully adaptive DPMM inference for high dimensional data. We propose a novel method that performs adaptive clustering with DPMM using collapsed VI, while incorporating weakly-informative priors for $\alpha$ and $G_{0}$. We illustrate the importance of $G_{0}$ covariance structure and prior choice by considering different parameterisations of the data covariance matrix. On high-dimensional Gaussian simulations, our model demonstrates substantially faster convergence than a state-of-the-art MCMC splice sampler. We further evaluate performances on Negative Binomial simulations and conduct sensitivity analyses to assess robustness on realistic data conditions. Application to a publicly available leukemia transcriptomic data set comprising 72 samples and 2,194 gene expression successfully recovers every known sub-type, all while identifying additional gene expression-based sub-clusters with meaningful biological interpretation.
keywords:
Clustering, Bayesian Nonparametrics, Dirichlet process mixture model, Variational inference, Unstructured covariance, Concentration parameter, High-dimension, Gene expression data
issue: x
1 Introduction
Clustering is an exploratory data analysis technique, where the goal is to group the data based on measured or estimated inherent similarities (jain2010data). Since the early emergence of $K$-means in 1950s (steinhaus1956division; forgy1965cluster; mcqueen1967some; lloyd1982least), a large number of clustering methods have been deployed for diverse interdisciplinary applications that deal with high dimensional data sets (dinh2025data). These methods include density based approaches (ester1996density; campello2013density), graph based techniques (ertoz2002new; traag2019louvain) and probabilistic models that define clusters based on mixture distributions (bock1996probabilistic, bouveyron2007high). However, a prerequisite for standard clustering techniques and finite mixture models is the number of clusters $K$ (or analogous parameters for density based methods) present a priori (cai2021finite). Even dimensionality reduction techniques like PCA (pearson1901liii) or UMAP (mcinnes2018umap) fail to capture underlying structure for high-dimensional data and require strong domain-specific knowledge to select an appropriate $K$ (wani2025comprehensive). To circumvent this limitation, Bayesian non-parametric mixture models (NPMM) have emerged as an efficient alternative to perform adaptive clustering based on probability distributions without pre-specifying $K$ (orbanz2010bayesian; gershman2012tutorial).
Introduction of Dirichlet process $DP(\alpha,G_{0})$ by ferguson1973bayesian enabled the development of DP mixture models (DPMM) as a class of Bayesian NPMM for density estimation, clustering as well as regression (muller2013bayesian). DPMM assume an infinite choice of unknown $K$ corresponding to partitions of the mixture components based on a DP prior (fruhwirth2019here), and corresponding posterior inference facilitates the estimate of $K$ (a non-empty partition corresponds to a cluster $k\in K$). Although DPMM have been scrutinized for inconsistent posterior estimation (miller2014inconsistency; miller2013simple), ascolani2023clustering have successfully demonstrated that consistency can be obtained by varying the concentration parameter $\alpha$, for instance, using a Gamma prior to implement a fully Bayesian scheme to update $\alpha$ (west1992hyperparameter).
DPMM often results in analytically intractable likelihoods and posteriors. Markov Chain Monte Carlo (MCMC) methods are a popular choice to sample from such analytically intractable distributions for parameter inference (escobar1995bayesian; neal2000markov). However, these methods exhibit slow convergence rates with difficulty in assessment (gershman2012tutorial), especially for high dimensions with several hundreds or even thousands of features per sample and limited sample size. An alternative is to use variational inference (VI) that approximates the true posterior $p(\cdot|data)$ with a tractable density function $q(\cdot)$ from a (known) family of probability distributions while deterministically optimising an information-theoretic criterion like KL divergence (blei2017variational). Rooted in early estimation methods for graphical models (jordan1999introduction), VI has gained prominence for its computational efficiency and scalability, particularly in modern machine learning applications involving large datasets (bernardi2024variational; loya2025scalable).
A general scheme of DPMM implementation would involve a stick-breaking representation of the DP prior (sethuraman1994constructive) with hyper-priors defined for $\alpha$ and base distribution $G_{0}$. To estimate model parameters, mean-field VI is the most widely adopted approach that assumes a factorized variational distribution and iteratively updates the parameters using conjugate priors (blei2006variational; bishop2006pattern). Although the independent factorization among variational parameters induces analytical tractability, latent allocation variables in DPMM are inherently dependent through their shared mixture proportions across samples. This fails to capture the relationships between hierarchical parameters and results in inconsistency between cluster numbers estimated from $\alpha$ and the latent allocation variables respectively (details in Supplementary material). kurihara2007collapsed have provided a collapsed VI framework by integrating out (collapsing) the mixture proportions, thus defining a prior distribution for the latent allocation vectors that depends only on $\alpha$ (1). This integration, however, constructs a non-standard prior with no standard conjugate hyper-prior for $\alpha$. Hence, the collapsed VI approach lacks a current working implementation with an adaptive inference on $\alpha$.
Apart from $\alpha$ that has a directly proportional effect on the inferred number of clusters (li2019tutorial), the model variance also plays a significant role in estimating $K$ (hejblum2019sequential). Incorporating $G_{0}$ variance as an unknown variable becomes important for coherent estimation, which is often ignored while implementing DPMM (including blei2006variational). The cluster variance particularly impacts high-dimensional data, which is often characterized by reduced signal-to-noise ratio, high variability due to limited sample size (johnstone2009statistical) and reduced separation between data points with increasing dimensions (clarke2008properties). The commonly used Inverse-Wishart $IW$ prior for covariance matrix receives criticism for a rigid structure, strongly coupled relations between variance and covariance (o2004kendall; tokuda2025visualizing) and high sensitivity to hyper-parameter choice (hennig2015handbook). Factorized covariance representations (like Cholesky decomposed factors) exhibit similar variance-covariance dependency as the resulting matrix have linear combination of constituent factor elements that are common in variance as well as covariance terms. Approaches like block diagonal covariance matrices, latent factor models (chandra2023escaping) and element-wise distributions (jing2024variance) focus on introducing conditional independence within the covariance matrices. A fundamental issue with Gaussian mixture models or GMM (both finite and infinite) is their assumption of Gaussian distributed clusters. This assumption is frequently violated in real world data sets, where cluster distributions may exhibit skewness, heavy tails, multimodality, or other deviations from normality (hejblum2019sequential; zhang2022bayesian). Paired with the ’curse of dimensionality’, data sets like gene expression data often show less separated or overlapped clusters (yu2017clustering) and using GMMs lead to misspecified clustering (kasa2023avoiding). In biological applications, the cluster overlapping can also arise from factors like multimodality and high correlation (clarke2008properties). In order to retrieve clustering estimates from such data, it is important to have cluster-specific distributions that regulate within-cluster spread as well as inter-cluster separability for robust outcomes.
In this paper, we propose a novel collapsed VI approach that explicitly incorporates the concentration parameter $\alpha$ as well as model covariance in a DPMM. Combining Euler-Maclaurin and Taylor series approximations, we obtain a closed form variational distribution for $\alpha$ with a conjugate Gamma hyper-prior. Following a selection between different covariance parameterisations, the analysis evaluates effect of hyper-prior choice through their corresponding performances on Gaussian simulated data. We establish the model with sparsity inducing hyper-prior (namely, Sparse DPMM) as the best choice (a desirable property for high-dimensional as well as highly-clustered data (jing2024variance)) and further evaluate the sensitivity based on the hyper-parameters with high dimensional ($d=1000$) Negative Binomial simulations. The optimal model is compared to an MCMC approach as implemented in the R package NPflow (hejblum2019sequential) in terms of convergence speed. We apply Sparse DPMM to cluster leukemia sub-types based on gene expression data set (armstrong2002mll; de2008clustering) with $2194$ genes and $72$ samples, and further evaluated the performance with existing clustering techniques. In addition, we provide a computational implementation through the R-package vimixr, available on CRAN.
The rest of the paper continues as follows. Section 2 develops the statistical methodology of our proposed model, focusing on formulating variational updates for different $G_{0}$ covariance hyper-priors and model implementation (Table 1). Section 3 highlights the theoretical merits and illustrates the inference results on simulated data for optimal model choice and sensitivity analysis with an emphasis on hyper-parameter tuning, while comparing computational performance with a splice sampling MCMC approach (hejblum2019sequential). It is followed by an application of Sparse DPMM on gene expression data to cluster acute lymphoblastic (ALL), mixed lineage (MLL) and acute myeloid (AML) leukemia sub-types (armstrong2002mll) and compare performance with $K$-means (mcqueen1967some), density-based DBSCAN (ester1996density) and HDBSCAN (campello2013density), graph-based shared nearest neighbourhood (ertoz2002new) and Leiden (kelly2023leiden), and model-based high dimensional clustering (bouveyron2007high) techniques. Section 4 concludes with discussions on the benefits and short comings of the proposed model, with possible extensions for further research and development.
2 Methods
2.1 Problem set-up
Consider $X_{n}\in\mathbb{R}^{d}$ random variables sampled from an unknown mixture of distributions $F$
$X_{n}|G\overset{\mathrm{i.i.d}}{\sim}F$ for $n=1,\dots,N$
where $F(\mathbf{X})=\int_{\Theta}f_{\eta}(\mathbf{X})G(d\eta)$ (hejblum2019sequential). We assume a multivariate Gaussian distribution for $f_{\eta}(.)$ with mean $\mu$ and covariance matrix $\Sigma$ ($\eta=\{\mu,\Sigma\}$). $G$ is the unknown mixing distribution that characterizes the mixture components over the parameter space $\Theta$, thus inducing clustering among $X_{n}$. A DP prior on $G\sim DP(\alpha,G_{0})$ with concentration parameter $\alpha$ and base distribution $G_{0}$ (ferguson1973bayesian) gives us a non-parametric mixing distribution
$G(.)=\sum_{k=1}^{\infty}\pi_{k}\delta_{\eta_{k}}(.)$ where $\eta_{k}\in G_{0}$
The mixing proportions $\pi_{k}$’s are drawn from a stick-breaking scheme (sethuraman1994constructive), which further introduce latent allocation variables $Z_{n}$ in the model that follows a categorical distribution with parameters $\{\pi_{k}\}$’s such that $p(Z_{n}=k)=\pi_{k}$ for every $n\in N$. The likelihood of $X_{n}$ hence follows
$X_{n}\sim\prod_{k}MVN(\mu_{k},\Sigma_{k})^{\mathbb{I}[Z_{n}=k]}$
(blei2006variational). Our objective is to estimate $Z_{n}$ as well as $\eta_{k}=\{\mu_{k},\Sigma_{k}\}$ to quantitatively evaluate the partitioning of the parameter space $\mathbf{\Theta}$.
2.2 Dirichlet process mixture model
A hierarchical Bayesian model follows
$V_{k}\sim Beta(1,\alpha)$
$\pi_{k}:=V_{k}\prod_{j<k}(1-V_{j})$
$\mu_{k},\Sigma_{k}\sim G_{0}$
$z_{n}\sim Categorical\left(\{\pi_{k}\}\right)$
$x_{n}|z_{n}\sim MVN(\mu_{z_{n}},\Sigma_{z_{n}})$
This is an illustration of a Dirichlet process mixture model (DPMM), introduced by antoniak1974mixtures. The concentration parameter $\alpha$ has a direct impact on the posterior expectation of the number of non-empty clusters (teh2010dirichlet). escobar1995bayesian have proposed a data augmentation scheme that leverages a Gamma hyper-prior $\alpha\sim Gamma(a,b)$ yielding a posterior distribution that adjust to the actual number of clusters observed in the data with consistency (ascolani2023clustering). With a conjugate choice for $G_{0}$, the model provides an estimate for the unknown parameters of interest.
2.3 Choice of $G_{0}$
The underlying structure of $\eta=\{\mu,\Sigma\}$ governs the distributional choice for $G_{0}$. We use a multivariate Gaussian $MVN(\mathbf{0},\Sigma_{\mu})$ as the prior for $\mu_{k}$. To illustrate the effect of variance parameter on a clustering model, several parameterisation for $\Sigma$ are presented in Table 1.
Assumption Structure Prior
Fixed $M_{1}:\frac{1}{\sigma}I_{d\times d}$
$M_{2}:\Sigma$
Unknown (global) $M_{3}:\frac{1}{\sigma}I_{d\times d}$ $\sigma\sim\Gamma(g_{1},g_{2})$
$M_{4}:\Sigma$ $\Sigma\sim IW(\nu_{0},V_{0})$
$M_{5}:\Sigma^{-1}=LL^{t}$ $L_{ij}\sim N(\mu_{0},\sigma_{0}),L_{ii}^{2}\sim\Gamma(a_{0},b_{0})$
Unknown (cluster-specific) $M_{6}:\Sigma_{k}$ $\Sigma_{k}\sim IW(\nu_{0},V_{0})$
$M_{7}:\Sigma_{k}$ $\Sigma_{k_{ij}}^{-1}\sim Lap(0,c_{0}),\Sigma_{k_{ii}}^{-1}\sim\Gamma(a_{0},b_{0})$
$M_{8}:\Sigma_{k}$ $\Sigma_{k_{ij}}^{-1}\sim N(c_{0},10^{-6}),\Sigma_{k_{ii}}^{-1}\sim\Gamma(a_{0},b_{0})$
Table 1: Model structure based on choice of $\Sigma$; $\sigma$ is a scalar quantity, $L$ is the Cholesky factorized lower triangular matrix for $\Sigma^{-1}=LL^{t}$ and $i,j\in{1,\dots,d}$
We introduce gradual complexity in terms of covariance matrix choices, starting from known or fixed variance ($M_{1},M_{2}$). For unknown variance, global refers to a common covariance matrix $\Sigma$ for all cluster choices ($M_{3},M_{4},M_{5}$), whereas cluster-specific defines each cluster with a unique covariance matrix $\Sigma_{k}$, along with the mean vectors ($M_{6},M_{7},M_{8}$). The choice of prior distributions maintain the conjugacy of the models, while considering hyper-parameters that establish weakly informative priors.
2.4 Posterior estimation using collapsed VI
We implement a collapsed variational inference (CVI) approach, where we integrate out (or collapse) the stick-breaking parameters $\{V_{k}\}$, following kurihara2007collapsed. This yields a distribution of the latent allocation vectors $\{z_{n}\}$ depending only on the DP concentration parameter $\alpha$:
$\begin{split}p(z_{n}|\alpha)&=\prod_{k}\int_{V_{k}}p(z_{n}|V_{k})p(V_{k}|\alpha)dV_{k}\\ &=\prod_{k}\alpha\frac{\Gamma(\mathbf{1}[z_{n}=k]+1)\Gamma(\mathbf{1}[z_{n}>k]+\alpha)}{\Gamma(1+\mathbf{1}[z_{n}\geq k]+\alpha)}\end{split}$ (1)
where $\mathbf{1}[.]$ represents the indicator function.
We can reconstruct our DPMM as:
$\alpha\sim Gamma(a,b)$
$\mu_{k},\Sigma_{k}\sim G_{0}$
$z_{n}\sim p(z_{n}|\alpha)$ (1)
$x_{n}|z_{n}\sim MVN(\mu_{z_{n}},\Sigma_{z_{n}})$.
The posterior distribution is approximated by a mean-field variational distribution for the DPMM parameters
$q(\alpha,\{\mu_{k}\},\{\Sigma_{k}\},\{z_{n}\})=q(\alpha)\prod_{k\leq K}q(\mu_{k})q(\Sigma_{k})\prod_{n}q(z_{n})$
(blei2017variational; zhang2018advances). Although having theoretically “infinite“ prior choices for $k$, truncated variational distributions with an upper limit $k=K$ enable practical implementation (ishwaran2001gibbs).
For posterior estimation, we use exponential families of distributions for $q(\cdot)$ and apply coordinate ascent algorithm to update the variational hyper-parameters (bishop2006pattern). It ensures guarantee of convergence and provides a closed form update $q^{*}(\cdot)$ that maximizes the Evidence lower bound (ELBO) (or equivalently minimizing the information-theoretic Kullbach-Leibler (KL) divergence)
$\begin{split}q^{*}(\theta_{i})&\propto\exp(\mathbb{E}_{q_{\theta^{-i}}}\left[\log\,p(X,\mathbf{\theta})\right])\\ &\propto\exp(\mathbb{E}_{q_{\theta^{-i}}}\left[\log\,p(X|\mathbf{\theta})+\log\,p(\mathbf{\theta})\right])\end{split}$ (2)
where $\theta_{j}\in\mathbf{\Theta}=\{\alpha,\{\mu_{k}\},\{\Sigma_{k}\},\{z_{n}\}\}$ and $\theta^{-j}=\mathbf{\Theta}\backslash\theta_{j}$ represents the set of all except the $j^{th}$ parameter (blei2006variational; bishop2006pattern). Appropriate prior as well as variational distribution choices facilitate the update of $q^{*}(\cdot)$ hyper-parameters. These updates are used to perform posterior inference for the unknown $\{z_{n}\},\{\theta_{i}\}$, thus estimating data cluster as well as distributional partitions corresponding to the cluster allocations.
2.4.1 Latent allocations $\{z_{n}\}$
After collapsing the stick-breaking parameters, the prior conditional distribution of $z_{n}$ can be written as
$p(z_{n}=k|\textbf{z}^{-n})=\frac{1+N_{k}^{-n}}{1+\alpha+N_{\geq k}^{-n}}\prod_{j<k}\frac{\alpha+N_{>j}^{-n}}{1+\alpha+N_{\geq j}^{-n}}$
(kurihara2007collapsed). So, the variational distribution updates as
$q^{*}(z_{n})\propto\exp\biggl(\sum_{k\leq K}\mathbf{1}[z_{n}=k]\mathbb{E}_{q^{-z_{n}}}\biggl[\log\,p(z_{n}=k|\textbf{z}^{-n})+log(p(x_{n}|\mu_{k},\Sigma_{k})\biggr]\biggr).$ (3)
$q^{*}(z_{n})$ follows a Categorical distribution with probability hyper-parameters $q_{nk}$:
$\begin{split}q_{nk}&\propto\exp\biggl[\log(1+\mathbb{E}_{q}[N_{k}^{-n}])-\frac{\mathbb{V}_{q}[N_{k}^{-n}]}{(1+\mathbb{E}_{q}[N_{k}^{-n}])^{2}}\\ &-\log(1+\mathbb{E}_{q}[N_{\geq k}^{-n}]+\frac{w_{1}}{w_{2}})+\frac{\mathbb{V}_{q}[N_{\geq k}^{-n}]+\frac{w_{1}}{w_{2}^{2}}}{(1+\mathbb{E}_{q}[N_{\geq k}^{-n}]+\frac{w_{1}}{w_{2}})^{2}}\\ &+\sum_{j<k}\left(\log(\frac{w_{1}}{w_{2}}+\mathbb{E}_{q}[N_{>j}^{-n}])-\frac{\mathbb{V}_{q}[N_{>j}^{-n}]+\frac{w_{1}}{w_{2}^{2}}}{(\frac{w_{1}}{w_{2}}+\mathbb{E}_{q}[N_{>j}^{-n}])^{2}}\right.\\ &\left.-\log(1+\frac{w_{1}}{w_{2}}+\mathbb{E}_{q}[N_{\geq j}^{-n}])+\frac{\mathbb{V}_{q}[N_{\geq j}^{-n}]+\frac{w_{1}}{w_{2}^{2}}}{(1+\frac{w_{1}}{w_{2}}+\mathbb{E}_{Q}[N_{\geq j}^{-n}])^{2}}\right)\\ &+\mathbb{E}_{q}[\log\,p(x_{n}|\mu_{k},\Sigma_{k})]\biggr]\end{split}$ (4)
where $N_{k}=\sum_{n}\mathbf{1}[z_{n}=k]$, $N_{>k}=\sum_{n}\mathbf{1}[z_{n}>k]$ and $N_{\geq k}=N_{k}+N_{>k}$. The calculations of their corresponding expectations and variances follow from kurihara2007collapsed (details in Supplementary material).
2.4.2 Distributional parameters $\{\mu_{k},\Sigma_{k}\}$
We apply multivariate Gaussian distributions, both as prior and variational $q(\cdot)$, for $\mu_{k}$ such that
$p(\mu_{k})\sim MVN(\mu_{0}=\mathbf{0},\Sigma_{\mu})$
$q(\mu_{k})\sim MVN(\phi_{k},\Lambda_{k})$.
The updated $q^{*}(\mu_{k})$ depends on $p(\mu_{k})$ as well as $p(X_{n}=x_{n}|z_{n})$, and hence the structure of $\Sigma_{k}$ plays an important role. For instance, when we assume $\Sigma_{k}$ to be fixed and consider $\Sigma_{k}=\Sigma=\frac{1}{\sigma}I_{d\times d}$, the variational update of $\mu_{k}$ follows blei2006variational. On the other hand, if we consider $\Sigma_{k}=\Sigma$ as a full matrix, the updated hyper-parameters of $\mu_{k}$ can be written as
$\begin{split}&\Lambda_{k}=\biggl(\Sigma_{\mu}^{-1}+(\sum_{n}q_{nk})\Sigma^{-1}\biggr)^{-1}\\ &\phi_{k}=\Lambda_{k}\Sigma^{-1}\biggl(\sum_{n}q_{nk}x_{n}\biggr)(\text{details in Supplementary material})\end{split}$ (5)
Analogous results are observed for unknown $\Sigma$ with global structures. Owing to the mean-field variational family, the expression of $\Sigma^{-1}$ in Eq:(5) is replaced by $\mathbb{E}_{q}[\Sigma^{-1}]$ with conjugate choices for $q(\Sigma)$ (details in Supplementary material).
For cluster-specific unknown $\Sigma_{k}$, the prior distribution of $\mu_{k}$ is conditionally dependent on $\Sigma_{k}$ so that $p(\mu_{k}|\Sigma_{k})\sim MVN(\mu_{0}=\mathbf{0},\frac{1}{k_{0}}\Sigma_{k})$, where $k_{0}$ is a fixed scaling factor and influences the scale of $\mu_{k}$’s. For full matrix structure, we apply $q(\Sigma_{k})\sim IW(\nu_{k},V_{k})$ and the hyper-parameters are updated as
$\begin{split}&\nu_{k}=\nu_{0}+1+\sum_{n}q_{nk}\\ &V_{k}=V_{0}+\sum_{n}q_{nk}x_{n}x_{n}^{t}\\ &\Lambda_{k}=\frac{\mathbb{E}_{q}[\Sigma_{k}^{-1}]^{-1}}{k_{0}+\sum_{n}q_{nk}}\\ &\phi_{k}=\frac{\sum_{n}q_{nk}x_{n}}{k_{0}+\sum_{n}q_{nk}}\end{split}$ (6)
An alternative is to use element wise distributions on the precision matrix $\Sigma_{k}^{-1}$ as mentioned in Table:1. A simplification of the identity $det(e^{\Sigma_{k}^{-1}})=e^{tr(\Sigma_{k}^{-1})}$ (details in Supplementary material) facilitates the use of Gamma distributions for diagonal elements of $\Sigma_{k}^{-1}$. If we consider Laplace distribution for the off-diagonal elements, we obtain the following update
$\begin{split}&p(\Sigma_{k_{ii}}^{-1})\sim\Gamma(a_{0},b_{0}),q(\Sigma_{k_{ii}}^{-1})\sim\Gamma(a_{k_{i}},b_{k_{i}})\\ &p(\Sigma_{k_{ij}}^{-1})\sim Lap(0,c_{0}),q(\Sigma_{k_{ij}}^{-1})\sim Lap(0,c_{k_{ij}})\\ &a_{k_{i}}=a_{0}+\sum_{n}q_{nk}+1\\ &b_{k_{i}}=b_{0}+\frac{1}{2}\sum_{n}q_{nk}x_{ni}^{2}\\ &c_{k_{ij}}=\biggl(\frac{1}{c_{0}}+\frac{1}{2}\sum_{n}q_{nk}|x_{ni}x_{nj}|\biggr)^{-1}\end{split}$ (7)
and for Gaussian distributed off-diagonal elements, we get
$\begin{split}&p(\Sigma_{k_{ij}}^{-1})\sim N(c_{0},10^{-6}),q(\Sigma_{k_{ij}}^{-1})\sim N(c_{k_{ij}},10^{-6})\\ &c_{k_{ij}}=c_{0}-\frac{10^{-6}}{2}\sum_{n}q_{nk}x_{ni}x_{nj}\end{split}$ (8)
where $i,j\in\{1,2,...,d\}$ and $i>j$ (details in Supplementary material). Update of $\mu_{k}$ remains as shown in Eq:(6) for both the cases.
2.4.3 Concentration parameter $\alpha$
We use Gamma distributions as the prior and variational distribution for $\alpha$
$p(\alpha)\sim\Gamma(a,b)$
$q(\alpha)\sim\Gamma(w_{1},w_{2})$
so that
$q^{*}(\alpha)\propto\exp(\mathbb{E}_{q^{-\alpha}}[\log\,p(\alpha)]+\mathbb{E}_{q^{-\alpha}}[\log\,p(\textbf{Z}|\alpha)])$
where $\textbf{Z}=\{z_{n}\}$. Applying Taylor series approximation and Euler-Maclaurin formula, we obtain the estimates of $w_{1}$ and $w_{2}$ (details in Supplementary material).
$w_{1}$ $=a+t-1$ (9a)
$w_{2}$ $=b+\sum_{k<t}\biggl[\log(a_{0}+\mathbb{E}_{q}[N_{\geq k}])-\frac{\mathbb{V}_{q}[N_{\geq k}]}{(a_{0}+\mathbb{E}_{q}[N_{\geq k}])^{2}}-\log(a_{0}+\mathbb{E}_{q}[N_{>k}])$
$\quad+\frac{\mathbb{V}_{q}[N_{>k}]}{(a_{0}+\mathbb{E}_{q}[N_{>k}])^{2}}\biggr]+\log(a_{0}+\mathbb{E}_{q}[N_{t}])-\frac{\mathbb{V}_{q}[N_{t}]}{(a_{0}+\mathbb{E}_{q}[N_{t}])^{2}}-\log(a_{0}+1)$ (9b)
where $t$ is the position of last non-zero cluster.
2.5 Implementation
Starting with an initial choice of latent allocation probabilities $q^{\circ}_{nk}$, the DPMM iteratively updates VI parameters based on the coordinate ascent algorithm (Eq: (2)) until there is no significant change in ELBO (blei2006variational)
$\begin{split}\text{ELBO}&=\mathbb{E}_{q}\biggl[\log\,p(X,Z,\{\mu_{k},\Sigma_{k}\},\alpha)\biggr]-\mathbb{E}_{q}\biggl[\log\,q(Z,\{\mu_{k},\Sigma_{k}\},\alpha)\biggr]\\ &=\mathbb{E}_{q}\biggl[\log\,p(X|Z,\{\mu_{k},\Sigma_{k}\},\alpha)+\log\,p(Z,\{\mu_{k},\Sigma_{k}\},\alpha)\biggr]\\ &-\mathbb{E}_{q}\biggl[\log\,q(Z,\{\mu_{k},\Sigma_{k}\},\alpha)\biggr].\end{split}$ (10)
ELBO acts as a lower bound of log evidence $logp(X)$ which minimizes reverse KL divergence between the target and variational distributions. The objective of VI is to update the variational parameters such that ELBO is maximized, thus approximating $p(\cdot)$ with $q(\cdot)$.
Due to non-convexity of the ELBO, VI is sensitive to the choice of initial parameters (blei2017variational). We circumvent this by choosing random initialization, and select the optimum starting values based on variational log likelihood (VLL) term $\mathbb{E}_{q}[\log\,p(X|Z,\{\mu_{k},\Sigma_{k}\},\alpha)](=\mathbb{E}_{q}[\log\,p(X|Z)])$ from Eq: (10). Although ELBO explicitly regularizes model complexity by including negative entropy of $q(\cdot)$ via $\mathbb{E}_{q}[\log\,q(Z,\{\mu_{k},\Sigma_{k}\},\alpha)]$, it is effected numerically especially for higher dimensions and can over-influence the ELBO values (details in Supplementary material). The VLL, on the other hand, implicitly handles model complexity through expectation over $q(\cdot)$, while providing a numerical fit of the data based on updated variational distribution. Thus, it makes sense to select the initial VI parameters that maximizes VLL.
3 Results
The proposed DPMM integrates a hierarchical approach in terms of varying the DP concentration parameter $\alpha$ in a collapsed VI method. The variational Gamma hyper-parameter $w_{1}$ depends on the number of non-zero clusters $t$ present in Z (9a), instead of total variational clusters $K$. For a Mean field VI (blei2006variational), the distribution of Z depends on $\alpha$ via the mixing proportions $\{\pi_{k}\}$’s, which can be non-zero for empty clusters (smaller proportions for unlikely clusters). Owing to the hierarchical dependency between $\alpha$ and $\{\pi_{k}\}$’s, the variational update of $\alpha$ thus considers all the plausible clusters based on non-zero $\{\pi_{k}\}$’s (which is often equal to $K$), rather than only the non-zero clusters $t$. This leads to an inconsistency between the number of clusters estimated between $\alpha$ and $Z$ (details in Supplementary material). However, for a collapsed VI, the distribution of Z directly depends on $\alpha$ (1). And with the theoretical approximations (2.4.3), our approach is able to maintain consistent estimates of cluster number through the variational updates of $\alpha$ hyper-parameters (9a), (9b).
The cluster estimates also depend on the structure as well as prior choice of $G_{0}$ variance (hejblum2019sequential). A common covariance structure (models $M_{3},M_{4}$ and $M_{5}$ from Table 1) can prove inadequate to recover clusters with heterogeneous covariances (biernacki2002assessing). Alternatively, a cluster-specific covariance structure (models $M_{6},M_{7}$ and $M_{8}$) is favoured, which integrates within-cluster variance in the model and complements the effect of $\alpha$ (Fig: 1). For cluster-specific covariance structure, an Inverse-Wishart prior often incorporates rigidity in the model ($M_{6}$) (o2004kendall; tokuda2025visualizing). We introduced independence between variance and covariance terms by assigning element-wise prior distributions (models $M_{7}$ and $M_{8}$) and applied conjugate $\Gamma(.)$ distributions for the diagonal elements. For the off-diagonal elements, we considered Gaussian distributions with arbitrarily small fixed variance (model $M_{8}$) and 0-centered Laplace distribution (model $M_{7}$).
3.1 Effect of $G_{0}$ Covariance Structure and Prior Choice
To compare the different model choices given in Table:1, we generated random $N=100$ multivariate Gaussian variables of dimension $d=2$ and grouped them into $K_{true}=2$ clusters. Using our developed R-package vimixr, we ran the models over generated data and calculated posterior number of clusters $K_{post}$ as well as adjusted Rand index (ARI) scores for $100$ simulation runs.
Figure 1: $K_{post}$ and ARI scores for different model choice
In Figure 1, the fixed variance models $M_{1}$ with diagonal and $M_{2}$ with full covariance matrices, respectively, act as known parameters and set the baseline for performance comparison. For unknown variance, global covariance structure shows comparative performance with diagonal covariance model $M3$ and significantly poor performance with both full covariance model $M4$ and Cholesky-decomposed covariance model $M5$. $M3$ employs uncorrelated mixture distribution for all the variational clusters, thus reducing rigidity and variance-covariance coupling. Compared to global covariance, cluster-specific covariance structure provides better results.
For different prior choices of the cluster-specific covariance structure, we further illustrated the effect of dimensionality and higher clusters by increasing dimension to $d=100$ and calculating $K_{post}$ and ARI for different $K_{true}=\{4,6,8,10\}$.
Figure 2: ARI for different cluster-specific models when $K_{true}$ is varying
Element-wise prior models show better adaptability with varying cluster number than $IW$ prior model in a high-dimensional setting (Figure 2). $M_{7}$ and $M_{8}$ have comparable performances (with $M_{7}$ slightly better than $M_{8}$). This can be attributed to the fact that a Laplace distribution is equivalent to a marginalized Gaussian distribution with Exponential prior on the variance parameter. Owing to its sparsity inducing effect (jing2024variance), we call $M_{7}$ as the Sparse DPMM, and consider it to be the best possible choice among the models defined in Table 1.
3.2 Effect of hyper-parameters of cluster distributions $\{\mu_{k},\Sigma_{k}\}$
Real high-dimensional data suffers from the curse of dimensionality due to low signal-to-noise ratio and limited sample size. Sparse DPMM is equiped with hyper-parameters of $\{\mu_{k},\Sigma_{k}\}$ that can weed out features with weaker contributions, as well as increase the effective sample size, thus reducing model variability which is otherwise an issue for low-n-high-d settings. Due to standard procedure of normalising such data sets (xia2023statistical), we chose $\mu_{0}=\mathbf{0}$ as the prior hyper-parameter for $\mu_{k}$ and $a_{0}=b_{0}$ for the diagonal elements of $\Sigma_{k}^{-1}$ (2.4.2). Along with $a_{0}=b_{0}$, the scaling factor $k_{0}$ for $p(\mu_{k}|\Sigma_{k})\sim MVN(\mathbf{0},\frac{1}{k_{0}}\Sigma_{k}^{-1})$ (2.4.2) plays a significant role in defining the posterior cluster distributions via their variational updates (7), thus effecting posterior estimation of $K$.
We performed the sensitivity analysis with $N=100,d=1000$ Negative Binomial variables clustered into $K_{true}=3$ groups to illustrate the robustness of Sparse DPMM on non-Gaussian data with more than $2$ clusters. For $1000$ random initialisations, we compared the VLL values obtained for $a)$ different $a_{0}=b_{0}$ with fixed $k_{0}=1$ (Figure 3(a)) and $b)$ different $k_{0}$ with fixed $a_{0}=b_{0}=0.001$ (Figure 3(b)).
Figure 3: Sensitivity analysis of Sparse DPMM over the hyper-parameters 3(a) $a_{0}=b_{0}$ and 3(b) $k_{0}$ for Negative Binomial simulations with $K_{true}=3$
For Sparse DPMM, the derivatives of VLL with respect to $a_{0}=b_{0}$ and $k_{0}$ can be expressed as
$\frac{\delta}{\delta a_{0}}VLL$ $=\frac{1}{2}\sum_{k}\left(d\sum_{n}q_{nk}^{\circ}\Psi^{1}(a_{0}+\sum_{n}q_{nk}^{\circ})\right$.
$\left.-\sum_{i}\biggl[\frac{a_{0}\sum_{n}q_{nk}^{\circ}+(\frac{\sum_{n}q_{nk}^{\circ}x_{n_{i}}^{2}}{2})^{2}-\frac{\sum_{n}q_{nk}^{\circ}\sum_{n}q_{nk}^{\circ}x_{n_{i}}^{2}}{2}}{(a_{0}+\frac{1}{2}\sum_{n}q_{nk}^{\circ}x_{n_{i}}^{2})^{2}}\biggr]\right)$ (11a)
$\frac{\delta}{\delta k_{0}}VLL$ $=\sum_{k}\biggl[\frac{1}{(k_{0}+\sum_{n}q_{nk}^{\circ})^{2}}\left(d\sum_{n}q_{nk}^{\circ}-(2-\sum_{n}q_{nk}^{\circ})\sum_{i}(\sum_{n}q_{nk}^{\circ}x_{n_{i}})^{2}\right)\biggr]$ (11b)
where $\{q_{nk}^{\circ}\}$ are a random allocation probability values and $\Psi^{1}(\cdot)$ is the trigamma function (details in Supplementary material). Functional behaviour of $\frac{\delta}{\delta a_{0}}VLL$ relies on $\{q_{nk}^{\circ}\}$, and VLL can either attain maxima or converge to a stabilising value with increasing $a_{0}=b_{0}$ (Figure 3(a)). For both the conditions, we estimate $a_{0}$ for which either $\frac{\delta}{\delta a_{0}}VLL=0$ or the maximum curvature point of $\frac{\delta}{\delta a_{0}}VLL$ that corresponds to the stabilising convergence of VLL (details in Supplementary material). In either case, the estimated $a_{0}$ represents the optimal value after which there is no further gain in the log likelihood, hence no additional prior information can be incorporated with higher $a_{0}$. Thus, for a random choice of $\{q_{nk}^{\circ}\}$, $a_{0}$ estimation inherently puts an upper bound on the degree of information required for maximising the VLL. This enables the model to threshold dimensions that have stronger contributions to effect the posterior estimation, and subsequently the VLL for a given data.
On the other hand, behaviour of $\frac{\delta}{\delta k_{0}}VLL$ is converging, and the value of convergence depends on $\{q_{nk}^{\circ}\}$ (Figure 3(b)). $k_{0}$ acts as the equivalent sample size of the $\mu_{k}$ prior (murphy2007conjugate), and scales the variational hyper-parameters of $\mu_{k}$ in Eq: (6). So, a smaller value of $\frac{1}{k_{0}}$ reduces the variability of $\mu_{k}$ which is equivalent of having a larger data sample. As the variational updates of $\mu_{k}$ in Eq: (6) is a weighted average between cluster-specific samples and prior mean, a natural choice is to consider $k0=N+1$ where N is the data sample size.
3.3 Computational scalability
We implemented our model framework in an R-package vimixr with all the $8$ parameterisations given in Table 1. For Sparse DPMM, the computation scales with sample size $N$ as order $\mathcal{O}(Nlog(N))$ due to the allocation variable $\mathbf{Z}$ updates with conditional expectation on $z_{n}|\mathbf{z}^{-n}$. The computational dependence for dimension $d$ is of order $\mathcal{O}(dlog(d)log(log(d))$ (details in Supplementary material). Although higher dimensions achieve convergence with less number of iterations, the rate is slightly higher due to computation of cluster-specific $d$-dimensional precision matrices.
For a slice-sampling based MCMC approach, we used the function DPMGibbsN from NPflow R package (hejblum2019sequential) and compared our Sparse DPMM for $N=100,d=100$ Gaussian data. Even with $1000$ MCMC iterations, DPMGibbsN often fails to achieve convergence. Sparse DPMM, on the other hand, demonstrates an increase in computational time by a magnitude of two orders ($\sim 100$ times) during bench-marking (details in Supplementary material).
3.4 Cancer sub-type estimation
We applied the proposed clustering framework on the gene-expression data set from leukemia sub-type study by armstrong2002mll, previously used as a benchmark for comparative analysis of clustering algorithms de2008clustering. armstrong2002mll showed that leukemia with rearrangement of the MLL gene (mixed lineage leukemia gene, now renamed KMT2A) should be classified as a distinct clinical entity rather than common acute lymphoblastic leukemia. The chromosomal rearrangement of MLL creates fusion genes with various partner genes that influence lineage commitment (krivtsov2007mll). For example, certain MLL rearrangements result in overexpression of HOXA9 and PRG1 genes, which are characteristic markers of acute myeloid leukemia (AML) rather than ALL (armstrong2002mll). Recent studies have further revealed the lineage plasticity of MLL-rearranged leukemia, showing dynamic transitions between lymphoid and myeloid phenotypes (chen2022single; janssens2024mll).
Working on the Affymetrix data from armstrong2002mll, de2008clustering provides a filtered set of $d=2194$ expressed genes for clustering $N=72$ leukemia samples into ALL-MLL-AML sub-types (Fig 4(a)). For a strong choice of prior hyper-parameters corresponding to Section 3.2, we obtained $K_{post}=3$ with ARI score of $0.92$ (Figure 4(b)).
Figure 4: PCA projection on the first 22 principal components for (a) labelled Leukemia sub-types based on $2194$ genes and (b) Sparse DPMM cluster estimates obtained using strong hyper-parameters
Following our empirical Bayes approach in Section 3.2, we estimated $a_{0}=b_{0}=28.9076$, and along with $k_{0}=N+1=73$, we applied Sparse DPMM for random initial probability allocation values and chose the best performing model based on VLL. To check the sensitivity of Sparse DPMM with a weaker prior hyper-parameter, we used $a_{0}=b_{0}=10$ instead of $50$ and clustered the same data set. For this particular choice of $a_{0}=b_{0}$, we obtained $K_{post}=4$ with ARI score $0.86$. (Figure 5).
Figure 5: PCA projection on the first 22 principal components for Sparse DPMM cluster estimates obtained using weaker hyper-prarameters
Figure 6: Heatmap showing comparison between labelled sub-types and Sparse DPMM (weaker) cluster estimates with respect to specifically expressed genes for the labelled sub-types (armstrong2002mll) for all the $72$ samples (genes in bold, probe-id in italics)
Comparing with known sub-types in Figure 4(a), Sparse DPMM with weaker hyper-parameters identifies an additional $4^{th}$ cluster with one sample replacing a labelled AML sub-type (Figure 5). Analogous to (armstrong2002mll), we plot a heat-map based on the available genes (de2008clustering) that are specifically expressed for the known sub-types to compare the lineage of the $4^{th}$ estimated cluster (Figure 6). The heat-map reveals a mixed expression pattern for the 4th cluster. The cluster shows elevated expression of genes characteristic of ALL sub-types (CD22 and CD24), suggesting lymphoblastic features. It also displays moderate expression of genes typically overexpressed in MLL sub-types (PROML1, FLT3 and ADAM10). In contrast, genes predominantly expressed in AML sub-types (DF, CTSD and BSG) are not dominantly expressed in this cluster. However, this expression pattern is not clear-cut, with the 4th cluster displaying characteristics intermediate between ALL and MLL rather than fitting cleanly into a single category. This ambiguous profile is consistent with the documented lineage plasticity of MLL-rearranged leukemias, which can undergo dynamic transitions between lymphoid and myeloid phenotypes during disease progression and treatment (chen2022single; janssens2024mll). The identification of this 4th cluster by Sparse DPMM thus appears to capture this biological reality of MLL heterogeneity rather than representing a technical artifact.
3.4.1 Performance benchmarking with current state of the art clustering techniques
For high-dimensional clustering analysis, there are certain popular choice of working algorithms with varying implementation techniques. Methods like DBSCAN (ester1996density) and HDBSCAN (campello2013density) are based on density of the data distributed over dimensional space, where data points in the low-density region are identified as noise/outliers and form clusters using core points. Shared nearest neighbours based method (ertoz2002new) works by grouping data points with higher similarity of shared neighbours across a k-nearest neighbour graph. bouveyron2007high proposed a high dimensional model based clustering that implements dimensionality reduction for each cluster with parsimonious covariance structure and estimate the Gaussian mixture model parameters using EM iterations. Another class of popular methods include graph based algorithms, like Leiden (an upgrade of Louvain) (traag2019louvain) that constructs a k-nearest neighbour graph from the data and detects communities (clusters) by maximizing graph modularity. Classical $K$-means (mcqueen1967some), on the other hand, is a partitioning based method that minimizes the within-cluster variance by iteratively assigning data points to their nearest centroids.
For performance on the Leukemia data set, we implemented the methods in R and chose the key parameters involved with these methods based on internal validation metrics specific to each method (details in Supplementary material). Based on corresponding optimal parameters, these methods are compared with Sparse DPMM implementation with empirical Bayes hyper-parameters, and we validated their performances with run-time (in seconds), estimated cluster numbers and ajusted Rand index (ARI). Due to diversity in methodology of these techniques, we avoided any internal metric for comparison (due to possible bias for particular technique(s)) and focused on external validation metrics. Figure 7 shows that Sparse DPMM is significantly slower ($\sim 8.091$ seconds per iteration) in comparison to alternative methods. However, the performance metrics based on estimated number of clusters and ARI scores validate the (relative) superiority of Sparse DPMM.
Figure 7: Performance benchmarking of popular clustering methods against Sparse DPMM; based on $a)$ implementation run-time in seconds (iterations are indicated with mini-bars), $b)$ estimated number of clusters and $c)$ corresponding ARI scores
4 Discussion
We developed a Bayesian non-parametric mixture model using collapsed variational inference (VI) to perform unsupervised clustering. Our approach hierarchically incorporates the Dirichlet process (DP) concentration parameter $\alpha$ as a variable in the DP mixture model (DPMM) with Gaussian kernels. Based on the importance of adaptive estimation of $\alpha$ (ascolani2023clustering), the proposed framework enhances the consistency of collapsed VI model developed by kurihara2007collapsed. We also illustrated the effect of prior for covariance matrix, and concluded that a cluster-specific Sparse DPMM is the best choice. Implementing our theoretical work into an R-package vimixr, the user is allowed to choose the model and prior choice based on parameterisations given in Table 1.
The Sparse DPMM provides significantly faster convergence than an MCMC slice sampling approach (hejblum2019sequential). Although the convergence speed varies with the dimension $d$ due to cluster-specific covariance structure, the Sparse DPMM performs robustly well for a large range of samples $N$, dimensions $d$ and true clusters $K_{true}$ present in the data. Comparing with conventional clustering techniques, the iterative nature of Sparse DPMM, added to the model structure complexity, leads to a slower convergence than the conventional techniques considered. However, it provides better clustering estimates with empirical Bayes hyper-parameters using gene expression data for leukemia sub-types (armstrong2002mll). Even with weaker hyper-parameter values, Sparse DPMM estimates are consistent with gene expression profiles, hence providing meaningful sub-clusters of the already known sub-types.
A natural extension of Sparse DPMM is to generalize beyond Gaussian kernels for the DP base distribution parameter $G_{0}$, which becomes eminent for real-world data. For instance, overdispersed count data is better modelled using Negative Binomial distributions (fernandez2022comparison). So, incorporating Negative Binomial kernels can aid in avoiding model misspecification for clustering such data sets like gene expression data (anders2010differential). Another aspect is to introduce probabilistic feature selection variables that could be updated simultaneously (tadesse2005bayesian) or guide the clustering process itself (rouanet2024bayesian). This can provide insights on the features defining the estimated clusters and prove beneficial, both computationally and inferentially, for high-dimensional data sets.
As an optimisation technique, VI algorithms heavily depend on the initial choice of latent variables (latent allocation vectors $\{z_{n}\}$ in our case) for a given data. We advocate random initialisation for these parameters, and the choice is based on variational log likelihood (VLL) values. For VI implementation, alternative approaches (zhang2018advances) can be explored for further refinement. We have also shown that the hyper-parameters’ choice for $G_{0}$ significantly effect the cluster estimation. Based on the analytical properties of VLL, empirical Bayes estimates for the hyper-parameters prove to be efficient for robust estimation. Due to their implied significance over effective dimensions and sample size, it becomes important to consider hierarchical hyper-priors for the hyper-parameter $a_{0}=b_{0}$. However, the conjugate prior for both unknown shape and rate parameters of a Gamma distribution does not have a closed form (miller1980bayesian). So, incorporating an $a_{0}=b_{0}$ hyper-prior would add another layer of analytical complexity, along with increased computational cost due to additional model parameters.
The exchangeability of mixture components in DPMM introduces label switching between atoms as they are ordered sequentially through the stick-breaking construction (papaspiliopoulos2008retrospective; hastie2015sampling). Following kurihara2007collapsed, we addressed the issue by ordering the latent allocation probability updates such that the clusters are arranged in decreasing proportions. Alternatively, fruhwirth2011dealing shows identification methods for finite mixture model, that could be extrapolated to non-parametric mixtures.
5 Conflicts of interest
The authors declare that they have no competing interests.
6 Funding
This work is supported by funds from the Programmes et équipements prioritaires de recherche (PEPR) Santé Numérique SMATCH Task 2.N1, under the Agence nationale de la recherche (ANR), France.
7 Software and Data availability
Software in the form of R package vimixr with complete documentation is available on CRAN (https://cran.r-project.org/web/packages/vimixr/index.html). The Leukemia sub-type gene expression data is publicly available online at https://schlieplab.org/Static/Supplements/CompCancer/Affymetrix/armstrong-2002-v2/. The package implementation is provided in the Zenodo repository https://doi.org/10.5281/zenodo.18405688, along with metadata and results.
8 Supplementary Material
Supplementary material is available online at http://biostatistics.oxfordjournals.org.
Acknowledgments
The authors thank the anonymous reviewers for their valuable suggestions. This work was supported by a French government grant managed by the Agence Nationale de la Recherche (ANR) under the France 2030 program (reference SMATCH ANR-22-PESN-0003). For heavy computation, the study used Cluster Curta, a project by Mésocentre de Calcul Intensif Aquitain (MCIA), France (https://redmine.mcia.fr/projects/cluster-curta)