### v2

# Scalable Dirichlet Process Mixture Models with Unknown Concentration and Adaptive Covariance for High-Dimensional Clustering Applied to Leukemia Transcriptomics

Abstract
Dirichlet process mixture models (DPMM) provide a flexible framework for clustering complex biological data without pre-specifying the number of clusters. Although Markov Chain Monte Carlo (MCMC) methods have bridged the gap between theory and application for such models, they scale poorly to high-dimensional data (e.g. omics data) and suffer from slow and difficult convergence. Variational Inference (VI) represents a faster alternative, but lacks implementation without overly simplifying assumptions such as mean field independence and known covariance. We propose a novel methodology that performs scalable inference of DPMM for clustering using collapsed VI, ensuring coherent and data-driven estimation of the number of clusters in a fully Bayesian scheme. To accommodate high-dimensional data, we introduce cluster-specific adaptive covariance modeling with sparsity-inducing priors, improving flexibility while mitigating variance–covariance coupling and overparameterization. The resulting algorithm scales efficiently with dimension and sample size, substantially reducing computation time compared to state-of-the-art MCMC splice sampler. Realistic simulation studies under Gaussian and negative-binomial settings demonstrate accurate recovery of cluster structure and robust performance across varying signal-to-noise regimes, supported by sensitivity analyses. Application on a publicly available leukemia transcriptomic dataset comprising 72 samples and 2,194 gene expressions successfully recovers established subtypes while revealing biological plasticity of one mixed lineage leukemia outlier sample. The proposed framework provides a computationally efficient and theoretically coherent solution for high-dimensional Bayesian clustering with unknown concentration and adaptive covariance.
Keywords: Bayesian non-parametrics; Dirichlet process mixture model; Gene expression data; High-dimensional clustering; Variational inference.
1 Introduction
Clustering is an exploratory data analysis technique, where the goal is to group the data based on measured or estimated inherent similarities (Jain, 2010). Since the early emergence of $K$-means from 1956 onward (Steinhaus and others, 1956; MacQueen, 1967), clustering has become increasingly prevalent in the analysis of complex biological datasets to identify and characterize their underlying structure, and a large number of methods have been deployed to deal with clustering high dimensional datasets. These methods include density based approaches (Ester et al., 1996; Campello et al., 2013), graph based techniques (Ertoz et al., 2002; Traag et al., 2019) and probabilistic models that define clusters based on mixture distributions (Bock, 1996; Bouveyron et al., 2007). However, a prerequisite for standard clustering techniques and finite mixture models is the number of clusters $K$ (or analogous parameters for density as well as graph based methods) present a priori. Even dimensionality reduction techniques like PCA (Pearson, 1901) and UMAP (McInnes et al., 2018) often fail to capture underlying structure for high-dimensional data and require strong domain-specific knowledge to select an appropriate $K$ (Wani, 2025). To circumvent this limitation, Bayesian non-parametric mixture models (NPMM) have emerged as an efficient alternative to perform clustering based on probability distributions without pre-specifying $K$ (Gershman and Blei, 2012).
Ferguson (1973) introduced Dirichlet process $DP(\alpha,G_{0})$ that led to the development of DP mixture models (DPMM) as a class of Bayesian NPMM for density estimation, clustering and regression (Müller and Mitra, 2013). DPMM use an infinite dimensional parameter space over the partition space of the observations thanks to a DP prior, yielding a posterior distribution with varying $K$. Although DPMM have been scrutinized for inconsistent posterior estimation of $K$ ($K$ being the number of non-empty clusters) by Miller and Harrison (2013), Ascolani et al. (2023) successfully demonstrated that consistency is guaranteed as long as the concentration parameter $\alpha$ is not fixed but data-driven (Escobar and West, 1995).
DPMM often results in analytically intractable likelihoods and posteriors. Markov Chain Monte Carlo (MCMC) methods are a popular choice to sample from such distributions for parameter inference (Neal, 2000). However, these methods exhibit slow convergence rates with difficulty in assessment (Gershman and Blei, 2012), especially for high dimensions with several hundreds or even thousands of features per sample and limited sample size. An alternative is to use variational inference (VI) that approximates the true posterior $p(\cdot|data)$ with a tractable density function $q(\cdot)$ from a (known) family of probability distributions, while optimising an information-theoretic criterion like KL divergence (Blei et al., 2017). Rooted in early estimation methods for graphical models (Jordan et al., 1999), VI has gained prominence for its computational efficiency and scalability, particularly in modern machine learning applications involving large datasets (Bernardi et al., 2024; Loya et al., 2025).
A general scheme of DPMM implementation would involve a stick-breaking representation of the DP prior (Sethuraman, 1994) with hyper-priors defined for $\alpha$ and base distribution $G_{0}$. For estimation, mean-field VI is a widely adopted approach that assumes factorized variational distribution and iteratively updates the parameters using conjugate priors (Blei and Jordan, 2006; Bishop and Nasrabadi, 2006). Although the independent factorization induces analytical tractability, latent allocation variables in DPMM are inherently dependent through their shared mixture proportions across samples. This fails to capture the relationships between hierarchical parameters and results in inconsistency between cluster numbers estimated from $\alpha$ and the latent allocation variables respectively (details in Supplementary material). Kurihara et al. (2007) provided a collapsed VI framework by integrating out (collapsing) the stick-breaking proportions, thus defining a prior distribution for the latent allocation vectors that depends only on $\alpha$ (1). This integration, however, constructs a non-standard prior with no standard conjugate hyper-prior for $\alpha$. Hence, the collapsed VI approach lacks a current working implementation with an adaptive inference on $\alpha$.
Apart from $\alpha$ that has a directly proportional effect on the inferred number of clusters (Teh and others, 2010), the model variance also plays a significant role in estimating $K$ (Frühwirth-Schnatter and Pyne, 2010; Hejblum et al., 2019). Incorporating $G_{0}$ variance as an unknown variable becomes important for coherent estimation, which is often ignored while implementing DPMM (including Blei and Jordan (2006)). The cluster variance particularly impacts high-dimensional data, which is often characterized by reduced signal-to-noise ratio, high variability due to limited sample size, and reduced separation between data points with increasing dimensions (Clarke et al., 2008). Commonly used Inverse-Wishart $IW$ prior receives criticisms for a rigid and too informative structure and strongly coupled relations between variance and covariance (Tokuda et al., 2025). Factorized covariance representations (like Cholesky decomposed factors) exhibit similar variance-covariance dependency as the resulting matrix share factor elements by linear combinations. Alternative approaches are block diagonal covariance matrices, latent factor models and element-wise distributions (Jing et al., 2024) that focus on introducing conditional independence within the covariance matrices. Although widely used, a fundamental issue with Gaussian mixture models (GMM) is that cluster distributions may exhibit significant deviations from normality (Hejblum et al., 2019), particularly in real-world data. Paired with the ’curse of dimensionality’, datasets like transcriptomics data often show less separated or overlapped clusters (Yu et al., 2017) and using GMM lead to misspecified clustering. In biological applications, the cluster overlapping can also arise from factors like multimodality and high correlation (Clarke et al., 2008). In order to retrieve clustering estimates from such data, it is important to have cluster-specific distributions that regulate within-cluster spread as well as inter-cluster separability for robust outcomes.
In this paper, we propose a novel collapsed VI approach that adaptively incorporates the concentration parameter $\alpha$ as well as model covariance in a DPMM. Combining Euler-Maclaurin and Taylor series approximations, we obtain a closed form variational distribution for $\alpha$ with a conjugate Gamma hyper-prior. Following a selection between different covariance parameterizations (Table 1), we formulate corresponding variational updates and evaluate the effect of hyper-prior choice through their performances on Gaussian simulated data. We establish the model with sparsity inducing hyper-prior (namely, Sparse DPMM) as the best choice (a desirable property for high-dimensional as well as highly-clustered data (Jing et al., 2024)) and further evaluate hyper-parameter sensitivity with high dimensional Negative Binomial simulations. The optimal model is compared to an MCMC approach as implemented in the R package NPflow (Hejblum et al., 2019) in terms of convergence speed. We apply Sparse DPMM to cluster acute lymphoblastic (ALL), mixed lineage (MLL) and acute myeloid (AML) leukemia sub-types based on gene expression dataset (Armstrong et al., 2002; De Souto et al., 2008) with $2194$ genes and $72$ samples, and benchmark the performance with existing $K$-means (MacQueen, 1967), density-based DBSCAN (Ester et al., 1996) and HDBSCAN (Campello et al., 2013), graph-based shared nearest neighbourhood (Ertoz et al., 2002) and Leiden (Traag et al., 2019), and model-based high dimensional clustering (Bouveyron et al., 2007) techniques. Sparse DPMM shows better recovery of known clusters than the conventional methods, while identifying an additional sub-cluster that corresponds to the plastic nature of MLL sub-type. Finally, we conclude with a general discussion of the proposed model, with possible directions for future research. We also provide a computational implementation through the R-package vimixr, available on CRAN.
2 Methods
2.1 Problem set-up
Consider $X_{n}\in\mathbb{R}^{d}$ random variables sampled from an unknown mixture of distributions $F$
$X_{n}|G\overset{\mathrm{i.i.d}}{\sim}F$ for $n=1,\dots,N$
where $F(\mathbf{X})=\int_{\Theta}f_{\eta}(\mathbf{X})G(d\eta)$ (Hejblum et al., 2019). We assume a multivariate Gaussian distribution for $f_{\eta}(.)$ with mean $\mu$ and covariance matrix $\Sigma$ ($\eta=\{\mu,\Sigma\}$). $G$ is the unknown mixing distribution that characterizes the mixture components over the parameter space $\Theta$, thus inducing clustering among $X_{n}$. A DP prior on $G\sim DP(\alpha,G_{0})$ with concentration parameter $\alpha$ and base distribution $G_{0}$ (Ferguson, 1973) gives us a non-parametric mixing distribution
$G(.)=\sum_{k=1}^{\infty}\pi_{k}\delta_{\eta_{k}}(.)$ where $\eta_{k}\in G_{0}$
The mixing proportions $\pi_{k}$’s are drawn from a stick-breaking scheme (Sethuraman, 1994), which further introduce latent allocation variables $Z_{n}$ in the model that follows a categorical distribution with parameters $\{\pi_{k}\}$’s such that $p(Z_{n}=k)=\pi_{k}$ for every $n\in N$. The likelihood of $X_{n}$ hence follows
$X_{n}\sim\prod_{k}MVN(\mu_{k},\Sigma_{k})^{\mathbb{I}[Z_{n}=k]}$
(Blei and Jordan, 2006). Our objective is to estimate $Z_{n}$ as well as $\eta_{k}=\{\mu_{k},\Sigma_{k}\}$ to quantitatively evaluate the partitioning of the parameter space $\mathbf{\Theta}$.
2.2 Dirichlet process mixture model
A hierarchical Bayesian model follows
$V_{k}\sim Beta(1,\alpha)$
$\pi_{k}:=V_{k}\prod_{j<k}(1-V_{j})$
$\mu_{k},\Sigma_{k}\sim G_{0}$
$z_{n}\sim Categorical\left(\{\pi_{k}\}\right)$
$x_{n}|z_{n}\sim MVN(\mu_{z_{n}},\Sigma_{z_{n}})$
This is an illustration of a Dirichlet process mixture model (DPMM), introduced by Antoniak (1974). The concentration parameter $\alpha$ has a direct impact on the posterior expectation of the number of non-empty clusters (Teh and others, 2010). Escobar and West (1995) have proposed a data augmentation scheme that leverages a Gamma hyper-prior $\alpha\sim Gamma(a,b)$ yielding a posterior distribution that adjust to the actual number of clusters observed in the data with consistency (Ascolani et al., 2023). With a conjugate choice for $G_{0}$, the model provides an estimate for the unknown parameters of interest.
2.3 Choice of $G_{0}$
The underlying structure of $\eta=\{\mu,\Sigma\}$ governs the distributional choice for $G_{0}$. We use a multivariate Gaussian $MVN(\mathbf{0},\Sigma_{\mu})$ as the prior for $\mu_{k}$. To illustrate the effect of variance parameter on a clustering model, several parameterization for $\Sigma$ are presented in Table 1.
Assumption Structure Prior
Fixed $M_{1}:\frac{1}{\sigma}I_{d\times d}$
$M_{2}:\Sigma$
Unknown (global) $M_{3}:\frac{1}{\sigma}I_{d\times d}$ $\sigma\sim\Gamma(g_{1},g_{2})$
$M_{4}:\Sigma$ $\Sigma\sim IW(\nu_{0},V_{0})$
$M_{5}:\Sigma^{-1}=LL^{t}$ $L_{ij}\sim N(\mu_{0},\sigma_{0}),L_{ii}^{2}\sim\Gamma(a_{0},b_{0})$
Unknown (cluster-specific) $M_{6}:\Sigma_{k}$ $\Sigma_{k}\sim IW(\nu_{0},V_{0})$
$M_{7}:\Sigma_{k}$ $\Sigma_{k_{ij}}^{-1}\sim Lap(0,c_{0}),\Sigma_{k_{ii}}^{-1}\sim\Gamma(a_{0},b_{0})$
$M_{8}:\Sigma_{k}$ $\Sigma_{k_{ij}}^{-1}\sim N(c_{0},10^{-6}),\Sigma_{k_{ii}}^{-1}\sim\Gamma(a_{0},b_{0})$
Table 1: Model structure based on choice of $\Sigma$; $\sigma$ is a scalar quantity, $L$ is the Cholesky factorized lower triangular matrix for $\Sigma^{-1}=LL^{t}$ and $i,j\in{1,\dots,d}$
We introduce gradual complexity in terms of covariance matrix choices, starting from known or fixed variance ($M_{1},M_{2}$). For unknown variance, global refers to a common covariance matrix $\Sigma$ for all cluster choices ($M_{3},M_{4},M_{5}$), whereas cluster-specific defines each cluster with a unique covariance matrix $\Sigma_{k}$, along with the mean vectors ($M_{6},M_{7},M_{8}$). The choice of prior distributions maintains conjugacy, while considering hyper-parameters that establish weakly informative priors.
2.4 Posterior estimation using collapsed VI
We implement a collapsed variational inference (CVI) approach, where we integrate out (or collapse) the stick-breaking parameters $\{V_{k}\}$, following Kurihara et al. (2007). This yields a distribution of the latent allocation vectors $\{z_{n}\}$ depending only on the DP concentration parameter $\alpha$:
$\begin{split}p(z_{n}|\alpha)&=\prod_{k}\int_{V_{k}}p(z_{n}|V_{k})p(V_{k}|\alpha)dV_{k}\\ &=\prod_{k}\alpha\frac{\Gamma(\mathbf{1}[z_{n}=k]+1)\Gamma(\mathbf{1}[z_{n}>k]+\alpha)}{\Gamma(1+\mathbf{1}[z_{n}\geq k]+\alpha)}\end{split}$ (1)
where $\mathbf{1}[.]$ represents the indicator function.
We can reconstruct our DPMM as:
$\alpha\sim Gamma(a,b)$
$\mu_{k},\Sigma_{k}\sim G_{0}$
$z_{n}\sim p(z_{n}|\alpha)$ (1)
$x_{n}|z_{n}\sim MVN(\mu_{z_{n}},\Sigma_{z_{n}})$.
The posterior distribution is approximated by a mean-field variational distribution for the DPMM parameters
$q(\alpha,\{\mu_{k}\},\{\Sigma_{k}\},\{z_{n}\})=q(\alpha)\prod_{k\leq K}q(\mu_{k})q(\Sigma_{k})\prod_{n}q(z_{n})$
(Blei et al., 2017). Although having theoretically “infinite“ prior choices for $k$, truncated variational distributions with an upper limit $k=K$ enable practical implementation (Ishwaran and James, 2001).
For posterior estimation, we use exponential families of distributions for $q(\cdot)$ and apply coordinate ascent algorithm to update the variational hyper-parameters (Bishop and Nasrabadi, 2006). It ensures guarantee of convergence and provides a closed form update $q^{*}(\cdot)$ that maximizes the Evidence lower bound (ELBO) (or equivalently minimizing the information-theoretic Kullbach-Leibler (KL) divergence)
$\begin{split}q^{*}(\theta_{i})&\propto\exp(\mathbb{E}_{q_{\theta^{-i}}}\left[\log\,p(X,\mathbf{\theta})\right])\\ &\propto\exp(\mathbb{E}_{q_{\theta^{-i}}}\left[\log\,p(X|\mathbf{\theta})+\log\,p(\mathbf{\theta})\right])\end{split}$ (2)
where $\theta_{j}\in\mathbf{\Theta}=\{\alpha,\{\mu_{k}\},\{\Sigma_{k}\},\{z_{n}\}\}$ and $\theta^{-j}=\mathbf{\Theta}\backslash\theta_{j}$ represents the set of all except the $j^{th}$ parameter (Blei and Jordan, 2006; Bishop and Nasrabadi, 2006). Appropriate prior as well as variational distribution choices facilitate the update of $q^{*}(\cdot)$ hyper-parameters. These updates are used to perform posterior inference for the unknown $\{z_{n}\},\{\theta_{i}\}$, thus estimating data cluster as well as distributional partitions corresponding to the cluster allocations.
2.4.1 Latent allocations $\{z_{n}\}$
After collapsing the stick-breaking parameters, the prior conditional distribution of $z_{n}$ can be written as
$p(z_{n}=k|\textbf{z}^{-n})=\frac{1+N_{k}^{-n}}{1+\alpha+N_{\geq k}^{-n}}\prod_{j<k}\frac{\alpha+N_{>j}^{-n}}{1+\alpha+N_{\geq j}^{-n}}$
(Kurihara et al., 2007). So, the variational distribution updates as
$q^{*}(z_{n})\propto\exp\biggl(\sum_{k\leq K}\mathbf{1}[z_{n}=k]\mathbb{E}_{q^{-z_{n}}}\biggl[\log\,p(z_{n}=k|\textbf{z}^{-n})+log(p(x_{n}|\mu_{k},\Sigma_{k})\biggr]\biggr).$ (3)
$q^{*}(z_{n})$ follows a Categorical distribution with probability hyper-parameters $q_{nk}$:
$\begin{split}q_{nk}&\propto\exp\biggl[\log(1+\mathbb{E}_{q}[N_{k}^{-n}])-\frac{\mathbb{V}_{q}[N_{k}^{-n}]}{(1+\mathbb{E}_{q}[N_{k}^{-n}])^{2}}\\ &-\log(1+\mathbb{E}_{q}[N_{\geq k}^{-n}]+\frac{w_{1}}{w_{2}})+\frac{\mathbb{V}_{q}[N_{\geq k}^{-n}]+\frac{w_{1}}{w_{2}^{2}}}{(1+\mathbb{E}_{q}[N_{\geq k}^{-n}]+\frac{w_{1}}{w_{2}})^{2}}\\ &+\sum_{j<k}\left(\log(\frac{w_{1}}{w_{2}}+\mathbb{E}_{q}[N_{>j}^{-n}])-\frac{\mathbb{V}_{q}[N_{>j}^{-n}]+\frac{w_{1}}{w_{2}^{2}}}{(\frac{w_{1}}{w_{2}}+\mathbb{E}_{q}[N_{>j}^{-n}])^{2}}\right.\\ &\left.-\log(1+\frac{w_{1}}{w_{2}}+\mathbb{E}_{q}[N_{\geq j}^{-n}])+\frac{\mathbb{V}_{q}[N_{\geq j}^{-n}]+\frac{w_{1}}{w_{2}^{2}}}{(1+\frac{w_{1}}{w_{2}}+\mathbb{E}_{Q}[N_{\geq j}^{-n}])^{2}}\right)\\ &+\mathbb{E}_{q}[\log\,p(x_{n}|\mu_{k},\Sigma_{k})]\biggr]\end{split}$ (4)
where $N_{k}=\sum_{n}\mathbf{1}[z_{n}=k]$, $N_{>k}=\sum_{n}\mathbf{1}[z_{n}>k]$ and $N_{\geq k}=N_{k}+N_{>k}$. The calculations of their corresponding expectations and variances follow from Kurihara et al. (2007) (details in Supplementary material).
2.4.2 Distributional parameters $\{\mu_{k},\Sigma_{k}\}$
We apply multivariate Gaussian distributions, both as prior and variational $q(\cdot)$, for $\mu_{k}$ such that
$p(\mu_{k})\sim MVN(\mu_{0}=\mathbf{0},\Sigma_{\mu})$
$q(\mu_{k})\sim MVN(\phi_{k},\Lambda_{k})$.
The updated $q^{*}(\mu_{k})$ depends on $p(\mu_{k})$ as well as $p(X_{n}=x_{n}|z_{n})$, and hence the structure of $\Sigma_{k}$ plays an important role. For instance, when we assume $\Sigma_{k}$ to be fixed and consider $\Sigma_{k}=\Sigma=\frac{1}{\sigma}I_{d\times d}$, the variational update of $\mu_{k}$ follows Blei and Jordan (2006). On the other hand, if we consider $\Sigma_{k}=\Sigma$ as a full matrix, the updated hyper-parameters of $\mu_{k}$ can be written as
$\begin{split}&\Lambda_{k}=\biggl(\Sigma_{\mu}^{-1}+(\sum_{n}q_{nk})\Sigma^{-1}\biggr)^{-1}\\ &\phi_{k}=\Lambda_{k}\Sigma^{-1}\biggl(\sum_{n}q_{nk}x_{n}\biggr)(\text{details in Supplementary material})\end{split}$ (5)
Analogous results are observed for unknown $\Sigma$ with global structures. Owing to the mean-field variational family, the expression of $\Sigma^{-1}$ in Eq:(5) is replaced by $\mathbb{E}_{q}[\Sigma^{-1}]$ with conjugate choices for $q(\Sigma)$ (details in Supplementary material).
For cluster-specific unknown $\Sigma_{k}$, the prior distribution of $\mu_{k}$ is conditionally dependent on $\Sigma_{k}$ so that $p(\mu_{k}|\Sigma_{k})\sim MVN(\mu_{0}=\mathbf{0},\frac{1}{k_{0}}\Sigma_{k})$, where $k_{0}$ is a fixed scaling factor and influences the scale of $\mu_{k}$’s. For full matrix structure, we apply $q(\Sigma_{k})\sim IW(\nu_{k},V_{k})$ and the hyper-parameters are updated as
$\begin{split}&\nu_{k}=\nu_{0}+1+\sum_{n}q_{nk}\\ &V_{k}=V_{0}+\sum_{n}q_{nk}x_{n}x_{n}^{t}\\ &\Lambda_{k}=\frac{\mathbb{E}_{q}[\Sigma_{k}^{-1}]^{-1}}{k_{0}+\sum_{n}q_{nk}}\\ &\phi_{k}=\frac{\sum_{n}q_{nk}x_{n}}{k_{0}+\sum_{n}q_{nk}}\end{split}$ (6)
An alternative is to use element wise distributions on the precision matrix $\Sigma_{k}^{-1}$ as mentioned in Table:1. A simplification of the identity $det(e^{\Sigma_{k}^{-1}})=e^{tr(\Sigma_{k}^{-1})}$ (details in Supplementary material) facilitates the use of Gamma distributions for diagonal elements of $\Sigma_{k}^{-1}$. If we consider Laplace distribution for the off-diagonal elements, we obtain the following update
$\begin{split}&p(\Sigma_{k_{ii}}^{-1})\sim\Gamma(a_{0},b_{0}),q(\Sigma_{k_{ii}}^{-1})\sim\Gamma(a_{k_{i}},b_{k_{i}})\\ &p(\Sigma_{k_{ij}}^{-1})\sim Lap(0,c_{0}),q(\Sigma_{k_{ij}}^{-1})\sim Lap(0,c_{k_{ij}})\\ &a_{k_{i}}=a_{0}+\sum_{n}q_{nk}+1\\ &b_{k_{i}}=b_{0}+\frac{1}{2}\sum_{n}q_{nk}x_{ni}^{2}\\ &c_{k_{ij}}=\biggl(\frac{1}{c_{0}}+\frac{1}{2}\sum_{n}q_{nk}|x_{ni}x_{nj}|\biggr)^{-1}\end{split}$ (7)
and for Gaussian distributed off-diagonal elements, we get
$\begin{split}&p(\Sigma_{k_{ij}}^{-1})\sim N(c_{0},10^{-6}),q(\Sigma_{k_{ij}}^{-1})\sim N(c_{k_{ij}},10^{-6})\\ &c_{k_{ij}}=c_{0}-\frac{10^{-6}}{2}\sum_{n}q_{nk}x_{ni}x_{nj}\end{split}$ (8)
where $i,j\in\{1,2,...,d\}$ and $i>j$ (details in Supplementary material). Update of $\mu_{k}$ remains as shown in Eq:(6) for both the cases.
2.4.3 Concentration parameter $\alpha$
We use Gamma distributions as the prior and variational distribution for $\alpha$
$p(\alpha)\sim\Gamma(a,b)$
$q(\alpha)\sim\Gamma(w_{1},w_{2})$
so that
$q^{*}(\alpha)\propto\exp(\mathbb{E}_{q^{-\alpha}}[\log\,p(\alpha)]+\mathbb{E}_{q^{-\alpha}}[\log\,p(\textbf{Z}|\alpha)])$
where $\textbf{Z}=\{z_{n}\}$. Applying Taylor series approximation and Euler-Maclaurin formula, we obtain the estimates of $w_{1}$ and $w_{2}$ (details in Supplementary material).
$w_{1}$ $=a+t-1$ (9a)
$w_{2}$ $=b+\sum_{k<t}\biggl[\log(a_{0}+\mathbb{E}_{q}[N_{\geq k}])-\frac{\mathbb{V}_{q}[N_{\geq k}]}{(a_{0}+\mathbb{E}_{q}[N_{\geq k}])^{2}}-\log(a_{0}+\mathbb{E}_{q}[N_{>k}])$
$\quad+\frac{\mathbb{V}_{q}[N_{>k}]}{(a_{0}+\mathbb{E}_{q}[N_{>k}])^{2}}\biggr]+\log(a_{0}+\mathbb{E}_{q}[N_{t}])-\frac{\mathbb{V}_{q}[N_{t}]}{(a_{0}+\mathbb{E}_{q}[N_{t}])^{2}}-\log(a_{0}+1)$ (9b)
where $t$ is the position of last non-zero cluster.
2.5 Implementation
Starting with an initial choice of latent allocation probabilities $q^{\circ}_{nk}$, the DPMM iteratively updates VI parameters based on the coordinate ascent algorithm (Eq: (2)) until there is no significant change in ELBO (Blei and Jordan, 2006)
$\begin{split}\text{ELBO}&=\mathbb{E}_{q}\biggl[\log\,p(X,Z,\{\mu_{k},\Sigma_{k}\},\alpha)\biggr]-\mathbb{E}_{q}\biggl[\log\,q(Z,\{\mu_{k},\Sigma_{k}\},\alpha)\biggr]\\ &=\mathbb{E}_{q}\biggl[\log\,p(X|Z,\{\mu_{k},\Sigma_{k}\},\alpha)+\log\,p(Z,\{\mu_{k},\Sigma_{k}\},\alpha)\biggr]\\ &-\mathbb{E}_{q}\biggl[\log\,q(Z,\{\mu_{k},\Sigma_{k}\},\alpha)\biggr].\end{split}$ (10)
ELBO acts as a lower bound of log evidence $logp(X)$ which minimizes reverse KL divergence between the target and variational distributions. Hence, objective of VI is to update the variational parameters such that ELBO is maximized, thus approximating $p(\cdot)$ with $q(\cdot)$.
Due to non-convexity of the ELBO, VI is sensitive to the choice of initial parameters (Blei et al., 2017). We circumvent this by choosing random initialization, and select the optimum starting values based on variational log likelihood (VLL) term $\mathbb{E}_{q}[\log\,p(X|Z,\{\mu_{k},\Sigma_{k}\},\alpha)](=\mathbb{E}_{q}[\log\,p(X|Z)])$ from Eq: (10). Although ELBO explicitly regularizes model complexity by including negative entropy of $q(\cdot)$ via $\mathbb{E}_{q}[\log\,q(Z,\{\mu_{k},\Sigma_{k}\},\alpha)]$, it can over-influence the ELBO due to large number of parameters in high dimensions (details in Supplementary material). The VLL, on the other hand, implicitly handles model complexity through expectation over $q(\cdot)$, while providing a numerical fit of the data based on updated variational distribution.
3 Results
The proposed DPMM integrates a hierarchical approach in terms of varying the DP concentration parameter $\alpha$ in a collapsed VI method. The variational Gamma hyper-parameter $w_{1}$ depends on the number of non-zero clusters $t$ present in Z (9a), instead of total variational clusters $K$. For a Mean field VI (Blei and Jordan, 2006), the distribution of Z depends on $\alpha$ via the mixing proportions $\{\pi_{k}\}$’s, which can be non-zero for empty clusters (smaller proportions for unlikely clusters). Owing to the hierarchical dependency between $\alpha$ and $\{\pi_{k}\}$’s, the variational update of $\alpha$ thus considers all the plausible clusters based on non-zero $\{\pi_{k}\}$’s (which is often equal to $K$), rather than only the non-zero clusters $t$. This leads to an inconsistency between the number of clusters estimated between $\alpha$ and $Z$ (details in Supplementary material). However, for a collapsed VI, the distribution of Z directly depends on $\alpha$ (1). And with the theoretical approximations (2.4.3), our approach is able to maintain consistent estimates of $K$ through the variational updates of $\alpha$ hyper-parameters (9a), (9b).
The cluster estimates also depend on the structure as well as prior choice of $G_{0}$ variance (Hejblum et al., 2019). A common covariance structure (models $M_{3},M_{4}$ and $M_{5}$ from Table 1) can prove inadequate to recover clusters with heterogeneous covariances (Biernacki et al., 2002). Alternatively, a cluster-specific covariance structure (models $M_{6},M_{7}$ and $M_{8}$) is favoured, which integrates within-cluster variance in the model and complements the effect of $\alpha$ (Fig: 1). For cluster-specific covariance structure, an Inverse-Wishart prior often incorporates rigidity in the model ($M_{6}$) (Tokuda et al., 2025). We introduced independence between variance and covariance terms by assigning element-wise prior distributions (models $M_{7}$ and $M_{8}$) and applied conjugate $\Gamma(.)$ distributions for the diagonal elements. For the off-diagonal elements, we considered Gaussian distributions with arbitrarily small fixed variance (model $M_{8}$) and 0-centered Laplace distribution (model $M_{7}$).
3.1 Effect of $G_{0}$ Covariance Structure and Prior Choice
To compare the different model choices given in Table:1, we generated random $N=100$ multivariate Gaussian variables of dimension $d=2$ and grouped them into $K_{true}=2$ clusters. Using our developed R-package vimixr, we ran the models over generated data and calculated posterior number of clusters $K_{post}$ as well as adjusted Rand index (ARI) scores for $100$ simulation runs.
Figure 1: $K_{post}$ and ARI scores for different model choice
In Figure 1, the fixed variance models $M_{1}$ with diagonal and $M_{2}$ with full covariance matrices act as known parameters and set the baseline for performance comparison. For unknown variance, global covariance structure shows comparative performance with diagonal covariance $M3$ and significantly poor performance with both full covariance $M4$ and Cholesky-decomposed covariance $M5$. $M3$ employs uncorrelated mixture distribution for all the variational clusters, thus reducing rigidity and variance-covariance coupling. Compared to global covariance, cluster-specific covariance structure provides better results.
For different prior choices of the cluster-specific covariance structure, we further illustrated the effect of dimensionality and higher clusters by increasing dimension to $d=100$ and calculating $K_{post}$ and ARI for different $K_{true}=\{4,6,8,10\}$.
Figure 2: ARI for different cluster-specific models when $K_{true}$ is varying
Element-wise prior models show better adaptability with varying cluster number than $IW$ prior model in a high-dimensional setting (Figure 2). $M_{7}$ and $M_{8}$ have comparable performances (with $M_{7}$ slightly better than $M_{8}$). This can be attributed to the fact that a Laplace distribution is equivalent to a marginalized Gaussian distribution with Exponential prior on the variance parameter. Owing to its sparsity inducing effect (Jing et al., 2024), we call $M_{7}$ as the Sparse DPMM, and consider it to be the best possible choice among the models defined in Table 1.
Analogous to real-world datasets, we sampled Negative Binomial simulations to study the effect of $G_{0}$ hyper-parameters for Sparse DPMM. Analytical calculations provide data-dependent updates, which play the role of empirical Bayes estimators for those hyper-parameters (details in Supplementary materials).
3.2 Computational scalability
We implemented our model framework in an R-package vimixr with all the $8$ parameterizations given in Table 1. For Sparse DPMM, the computation scales with sample size $N$ as order $\mathcal{O}(Nlog(N))$ due to the allocation variable $\mathbf{Z}$ updates with conditional expectation on $z_{n}|\mathbf{z}^{-n}$. The computational dependence for dimension $d$ is of order $\mathcal{O}(dlog(d)log(log(d))$ (details in Supplementary material). Although higher dimensions achieve convergence with less number of iterations, the rate is slightly higher due to computation of cluster-specific $d$-dimensional precision matrices.
For a slice-sampling based MCMC approach, we used the function DPMGibbsN from NPflow R package (Hejblum et al., 2019) and compared our Sparse DPMM for $N=100,d=100$ Gaussian data. Even with $1000$ MCMC iterations, DPMGibbsN often fails to achieve convergence. Sparse DPMM, on the other hand, demonstrates an increase in computational time by a magnitude of two orders ($\sim 100$ times; details in Supplementary material).
3.3 Cancer sub-type estimation
We applied the proposed clustering framework on the gene-expression dataset from leukemia sub-type study by Armstrong et al. (2002), previously used as a benchmark for comparative analysis of clustering algorithms (De Souto et al., 2008). Armstrong et al. (2002) showed that leukemia with rearrangement of the MLL gene (mixed lineage leukemia gene, now renamed KMT2A) should be classified as a distinct clinical entity rather than common acute lymphoblastic leukemia. The chromosomal rearrangement of MLL creates fusion genes with various partner genes that influence lineage commitment (Krivtsov and Armstrong, 2007). For example, certain MLL rearrangements result in overexpression of HOXA9 and PRG1 genes, which are characteristic markers of acute myeloid leukemia (AML) rather than ALL (Armstrong et al., 2002). Recent studies have further revealed the lineage plasticity of MLL-rearranged leukemia, showing dynamic transitions between lymphoid and myeloid phenotypes (Chen et al., 2022; Janssens et al., 2024).
Working on the Affymetrix data from Armstrong et al. (2002), De Souto et al. (2008) provides a filtered set of $d=2194$ expressed genes for clustering $N=72$ leukemia samples into ALL-MLL-AML sub-types (Fig 3(a)). For a strong choice of prior hyper-parameters, we obtained $K_{post}=3$ with ARI score of $0.92$ (Figure 3(b)).
Figure 3: PCA projection on the first 22 principal components for (a) labelled Leukemia sub-types based on $2194$ genes and (b) Sparse DPMM cluster estimates obtained using strong hyper-parameters
Following our empirical Bayes approach, we estimated $a_{0}=b_{0}=28.9076$, and along with $k_{0}=N+1=73$, we applied Sparse DPMM for random initial probability allocation values and chose the best performing model based on VLL. To check the sensitivity of Sparse DPMM with a weaker prior hyper-parameter, we used $a_{0}=b_{0}=10$ instead of $50$ and clustered the same dataset. For this particular choice of $a_{0}=b_{0}$, we obtained $K_{post}=4$ with ARI score $0.86$. (Figure 4(a)).
(a) Sparse DPMM clusters using weaker hyper-prarameters
(b) Heat map of clusters based on sub-type specific expressed genes
Figure 4: 4(a)) PCA projection on the first 22 principal components for Sparse DPMM clusters obtained using weaker hyper-prarameters; 4(b)) Heatmap showing comparison between labelled sub-types and Sparse DPMM (weaker) cluster estimates with respect to specifically expressed genes for the labelled sub-types (Armstrong et al., 2002) for all the $72$ samples (genes in bold, probe-id in italics)
Comparing with known sub-types in Figure 3(a), Sparse DPMM with weaker hyper-parameters identifies an additional $4^{th}$ cluster with one sample replacing a labelled AML sub-type (Figure 4(a)). Analogous to (Armstrong et al., 2002), we plot a heat-map based on the available genes (De Souto et al., 2008) that are specifically expressed for the known sub-types to compare the lineage of the $4^{th}$ estimated cluster (Figure 4(b)). The heat-map reveals a mixed expression pattern for the 4th cluster. The cluster shows elevated expression of genes characteristic of ALL sub-types (CD22 and CD24), suggesting lymphoblastic features. It also displays moderate expression of genes typically overexpressed in MLL sub-types (PROML1, FLT3 and ADAM10). In contrast, genes predominantly expressed in AML sub-types (DF, CTSD and BSG) are not dominantly expressed in this cluster. However, this expression pattern is not clear-cut, with the 4th cluster displaying characteristics intermediate between ALL and MLL rather than fitting cleanly into a single category. This ambiguous profile is consistent with the documented lineage plasticity of MLL-rearranged leukemias, which can undergo dynamic transitions between lymphoid and myeloid phenotypes during disease progression and treatment (Chen et al., 2022; Janssens et al., 2024). The identification of this 4th cluster by Sparse DPMM thus appears to capture this biological reality of MLL heterogeneity rather than representing a technical artifact.
3.3.1 Performance benchmarking with current state of the art clustering techniques
For high-dimensional clustering analysis, there are certain popular choice of working algorithms with varying implementation techniques. Methods like DBSCAN (Ester et al., 1996) and HDBSCAN (Campello et al., 2013) are based on density of the data distributed over dimensional space, where data points in the low-density region are identified as noise/outliers and form clusters using core points. Shared nearest neighbours based method (Ertoz et al., 2002) works by grouping data points with higher similarity of shared neighbours across a k-nearest neighbour graph. Bouveyron et al. (2007) proposed a high dimensional model based clustering that implements dimensionality reduction for each cluster with parsimonious covariance structure and estimate the Gaussian mixture model parameters using EM iterations. Another class of popular methods include graph based algorithms, like Leiden (an upgrade of Louvain, Traag et al. (2019)) that constructs a k-nearest neighbour graph from the data and detects communities (clusters) by maximizing graph modularity. Classical $K$-means (MacQueen, 1967), on the other hand, is a partitioning based method that minimizes the within-cluster variance by iteratively assigning data points to their nearest centroids.
Figure 5: Performance benchmarking of popular clustering methods against Sparse DPMM; based on $a)$ implementation run-time in seconds (iterations are indicated with mini-bars), $b)$ estimated number of clusters and $c)$ corresponding ARI scores
For performance on the Leukemia dataset, we implemented the methods in R and chose the key parameters involved with these methods based on internal validation metrics specific to each method (details in Supplementary material). Based on corresponding optimal parameters, these methods are compared with Sparse DPMM implementation with empirical Bayes hyper-parameters, and we validated their performances with run-time (in seconds), estimated cluster numbers and ajusted Rand index (ARI). Due to diversity in methodology of these techniques, we avoided any internal metric for comparison across methods and focused on external validation metrics. Figure 5 shows that Sparse DPMM is significantly slower ($\sim 8.091$ seconds per iteration) in comparison to alternative methods. However, the performance metrics based on estimated number of clusters and ARI scores validate the (relative) superiority of Sparse DPMM.
4 Discussion
We developed a Bayesian non-parametric mixture model using collapsed variational inference (VI) to perform unsupervised clustering. Our approach hierarchically incorporates the Dirichlet process (DP) concentration parameter $\alpha$ as a variable in the DP mixture model (DPMM) with Gaussian kernels. Based on the importance of adaptive estimation of $\alpha$ (Ascolani et al., 2023), the proposed framework enhances the consistency of collapsed VI model developed by Kurihara et al. (2007). We also illustrated the effect of prior for covariance matrix, and concluded that a cluster-specific Sparse DPMM is the best choice. Implementing our theoretical work into an R-package vimixr, we provide a probabilistic tool to perform exploratory analysis of unstructured and high-dimensional biological data with significant potential applicability.
The Sparse DPMM provides significantly faster convergence than an MCMC slice sampling approach (Hejblum et al., 2019). Comparing with conventional clustering techniques, the iterative nature of Sparse DPMM, added to the model structure complexity, leads to a slower convergence. However, it provides better clustering estimates with empirical Bayes hyper-parameters using gene expression data for leukemia sub-types (Armstrong et al., 2002). Even with weaker hyper-parameters, Sparse DPMM estimates are consistent with gene expression profiles, hence providing a biologically meaningful sub-cluster of the known sub-types.
A natural extension of Sparse DPMM is to generalize beyond Gaussian kernels for $G_{0}$, which becomes eminent for real-world data. For instance, overdispersed count data is better modelled using Negative Binomial distributions, thus avoiding model misspecification for clustering such datasets like gene expression data (Anders and Huber, 2010). Another aspect is to introduce probabilistic feature selection variables that could be updated simultaneously (Tadesse et al., 2005) or guide the clustering process itself (Rouanet et al., 2024). This can provide insights on the features defining the estimated clusters and prove beneficial for high-dimensional datasets.
As an optimization technique, VI algorithms heavily depend on the initial choice of latent variables (latent allocation vectors $\{z_{n}\}$ in our case) for a given data. We advocate random initialization for these parameters, and the choice is based on variational log likelihood (VLL) values. For VI implementation, alternative approaches (Blei et al., 2017) can be explored for further refinement. Based on the analytical properties of VLL, empirical Bayes estimates for the $G_{0}$ hyper-parameters prove to be efficient for robust estimation. Due to their implied significance over effective dimensions and sample size, it becomes important to consider hierarchical hyper-priors for the hyper-parameter $a_{0}=b_{0}$. However, the conjugate prior for both unknown shape and rate parameters of a Gamma distribution does not have a closed form (Miller, 1980). So, incorporating an $a_{0}=b_{0}$ hyper-prior would add another layer of analytical complexity, along with increased computational cost due to additional model parameters.
The exchangeability of mixture components in DPMM introduces label switching between atoms as they are ordered sequentially through the stick-breaking construction (Papaspiliopoulos and Roberts, 2008). Following Kurihara et al. (2007), we addressed the issue by ordering the latent allocation probability updates such that the clusters are arranged in decreasing proportions. Alternatively, Frühwirth-Schnatter (2011) shows identification methods for finite mixture model, that can be extrapolated to non-parametric mixtures.
Acknowledgments
This work was supported by the University of Bordeaux’s Digital Public Health Graduate School, financed by the PIA 3 through the Agence Nationale de la Recherche (Ref: 17-EURE-0019), by the Programme et Equipement Prioritaire de Recheche Santé Numérique (PEPR SN) project SMATCH (Ref: 22-PESN-0003). Computer time for this study was provided by the computing facilities MCIA (Mésocentre de Calcul Intensif Aquitain) of the Université de Bordeaux and of the Université de Pau et des Pays de l’Adour.
Supplementary Material
Supplementary material is available with this paper at the Biometrics website on Wiley Online Library.
Appendix A
A.1 Software and Data availability
Software in the form of R package vimixr with complete documentation is available on CRAN (https://cran.r-project.org/web/packages/vimixr/index.html). The Leukemia sub-type gene expression data is publicly available online at
https://schlieplab.org/Static/Supplements/CompCancer/Affymetrix/armstrong-2002-v2/. The package implementation is provided in the Zenodo repository
https://doi.org/10.5281/zenodo.18405687, along with metadata and results.